OpenStudy (kise_ryouta):

1/5(25−5a)=4−a −1−1 0 no solution all real numbers

OpenStudy (benji):

Alright so we need to get rid of the 1/5, how do you undo fractions?

OpenStudy (kise_ryouta):

divide right

OpenStudy (benji):

If you think about it, a fraction is dividing already it is 1 divided by 5.

OpenStudy (benji):

We would multiply by the denominater to get rid of it so \[(5)\frac{ 1 }{ 5}(25-5a)=(5)(4-a)\]

OpenStudy (benji):

So the 6 and 1/5 cancel and we are left with \[25-5a=(5)(4-a)\]So what do we do next?

OpenStudy (kise_ryouta):

Combine like terms

OpenStudy (benji):

Oi what did you get

jhonyy9 (jhonyy9):

@Benji nice to meet you dear my friend this is more much easy 1 --(25-5a) = 4 -a like an easy way just factorize out the 5 from parentheses 5 and simplify it with this 5 from denominator of this fraction and will get what ? the best easy way to solve this exercise

OpenStudy (benji):

You can factor out a 5 and you would have \[\frac{ 1 }{ 5 }5(5-a)=4-a\]and then\[5-a=4-a\]

jhonyy9 (jhonyy9):

exactly @Benji and in this way you can solve it for ,,a" very very easy